# LQG inspired spacetimes as solutions of the Einstein equations

### v1

@@ V Polymerized BH
Another spacetime inspired by LQG is the polymerized BH. This spacetime is regular and, like the previous cases, has a throat hidden by an event horizon, thus being a model of a BB. The line element that describes this spacetime is written as [86]:
@@
The constant $k>0$ is the polymerized parameter and represents the radius of the throat with a minimal area $4\pi k^{2}$. It is interesting that there is also an event horizon, independent of the value of $k$, different from the Simpson-Visser spacetime. The horizon is located at $r_{h}=\sqrt{k^{2}+4m^{2}}$, and the radius of the throat is consistently smaller than the radius of the event horizon in this coordinate system.
If we try to map this spacetime as a solution of the equations, magnetically charged, the source can be described as a phantom scalar field with a NED, described by the functions:
@@
$L_{F}(r)$ $=$ $-\frac{\sqrt{1-\frac{k^{2}}{r^{2}}}\left(3k^{2}+2r^{2}\right)}{4q^{2}}+\frac{3k^{2}m+r^{3}}{2q^{2}r},\quad F(r)=\frac{2q^{2}}{r^{4}}.$ (46)
@@
[…]
In the week field limit, we find $L(F)\approx 32/15k^{2}-Fk^{2}/2q^{2}$.
@@
$L(r)$ $=$ $\frac{\sqrt{1-\frac{k^{2}}{r^{2}}}\left(27k^{4}+16k^{2}r^{2}+32r^{4}\right)}{15k^{2}r^{4}}-\frac{18k^{2}m}{5r^{5}},\quad L_{F}(r)=\frac{4q^{2}r^{2}\sqrt{1-\frac{k^{2}}{r^{2}}}}{3k^{4}+2r\sqrt{1-\frac{k^{2}}{r^{2}}}\left(3k^{2}m+r^{3}\right)-k^{2}r^{2}-2r^{4}},$ (48)
$F(r)$ $=$ $-\frac{\left(-3k^{4}-2r\sqrt{1-\frac{k^{2}}{r^{2}}}\left(3k^{2}m+r^{3}\right)+k^{2}r^{2}+2r^{4}\right)^{2}}{8q^{2}r^{8}\left(1-\frac{k^{2}}{r^{2}}\right)}.$ (49)
@@
[…]
@@
Figure 1: Plot of the functions $F(r)$, $F(P)$, and $L(F)$ to the polymerized solution (43), with $q=k=0.5m$. The extreme points $P_{1}$ and $P_{2}$ inform us about the points where the electromagnetic Lagrangian changes its behavior.
@@ VI Energy Conditions
The energy conditions for the Simpson-Visser case have already been analyzed for each field in the literature [30]. Therefore, we now need to check for solutions inspired by LQG.
@@ VI.2 Polymerized solution
$DEC_{2}[\phi]=2WEC_{3}[\phi]\Longrightarrow\frac{28k^{2}m}{5r^{5}}-\frac{2\sqrt{1-\frac{k^{2}}{r^{2}}}\left(21k^{4}+8k^{2}r^{2}+16r^{4}\right)}{15k^{2}r^{4}}\geq 0.$ (73)
@@ VII Conclusion
The first model we studied was the Simpson-Visser spacetime. This model was chosen because it has been extensively studied in the literature and served as a consistency test for our method. The results obtained, both for magnetic and electric charges, differ from those presented in [30, 79]. However, this difference arises because of the chosen coordinate system, so despite being different, our results are consistent with those found in the literature.
@@
Once the fields have been separated, we analyzed the energy conditions for each field individually. Since the Simpson-Visser case has been extensively studied in the literature, we did not address it here. For the second model, we found that all the energy conditions related to the electromagnetic field are always satisfied, whereas the scalar field can violate all the energy conditions, both inside and outside the event horizon, which is consistent with a phantom scalar field. For the third model, some of the analytical expressions are not very clear. However, we observe that, for points far from the event horizon, the inequality related to $NEC[E]$ is not satisfied, ensuring that, at least in some regions, the NEC is violated. Moreover, for the scalar field, it becomes more evident that the NEC is violated in all regions of spacetime. In summary, NEC violations are consistent with the behavior expected for BB solutions, such as those discussed here [87].
The stability properties of solutions are known to depend critically on their source characteristics [88, 89, 90]. Consequently, classifying source types capable of producing regular solutions is essential, the principal aim of this preliminary investigation. In future work, we intend to analyze more characteristics of these solutions that are influenced by their types of sources, such as stability, thermodynamics, and shadows.

### v2

@@ V Polymerized BH
Another spacetime inspired by LQG is the polymerized BH. This spacetime is regular and, like the previous cases, has a throat hidden by an event horizon, thus being a model of a BB. The line element that describes this spacetime is written as [Peltola:2008pa]:
@@
The constant $k>0$ is the polymerized parameter and represents the radius of the throat with a minimal area $4\pi k^{2}$. It is interesting that there is also an event horizon, independent of the value of $k$, different from the Simpson-Visser spacetime. The horizon is located at $r_{h}=\sqrt{k^{2}+4m^{2}}$, and the radius of the throat is consistently smaller than the radius of the event horizon in this coordinate system. In addition to the original article [Peltola:2008pa], this polymerized solution was explored in the analysis of gravitational waves and nearly circular spiral orbits [Yang:2024cnd] as well as in the study of gravitational lensing effects and light deflection in the strong and weak field regimes [KumarWalia:2022ddq]. If we try to map this spacetime as a solution of the equations, magnetically charged, the source can be described as a phantom scalar field with a NED, described by the functions:
@@
$L_{F}(r)$ $=$ $-\frac{\sqrt{1-\frac{k^{2}}{r^{2}}}\left(3k^{2}+2r^{2}\right)}{4q_{m}^{2}}+\frac{3k^{2}m+r^{3}}{2q_{m}^{2}r},\quad F(r)=\frac{2q_{m}^{2}}{r^{4}}.$ (48)
@@
$L(F)=\frac{3\ 2^{3/4}F^{5/4}k^{2}m}{5q_{m}^{5/2}}+\frac{\sqrt{2}\sqrt{F}}{q_{m}}+\sqrt{2-\frac{\sqrt{2}\sqrt{F}k^{2}}{q_{m}}}\frac{\left(-9\sqrt{2}Fk^{4}-14\sqrt{F}k^{2}q_{m}+32\sqrt{2}q_{m}^{2}\right)}{30k^{2}q_{m}^{2}}.$ (49)
In the week field limit, we find $L(F)\approx 32/15k^{2}-Fk^{2}/2q_{m}^{2}$.
@@
$L(r)$ $=$ $\frac{\sqrt{1-\frac{k^{2}}{r^{2}}}\left(27k^{4}+16k^{2}r^{2}+32r^{4}\right)}{15k^{2}r^{4}}-\frac{18k^{2}m}{5r^{5}},\quad L_{F}(r)=\frac{4q_{e}^{2}r^{2}\sqrt{1-\frac{k^{2}}{r^{2}}}}{3k^{4}+2r\sqrt{1-\frac{k^{2}}{r^{2}}}\left(3k^{2}m+r^{3}\right)-k^{2}r^{2}-2r^{4}},$ (50)
$F(r)$ $=$ $-\frac{\left(-3k^{4}-2r\sqrt{1-\frac{k^{2}}{r^{2}}}\left(3k^{2}m+r^{3}\right)+k^{2}r^{2}+2r^{4}\right)^{2}}{8q_{e}^{2}r^{8}\left(1-\frac{k^{2}}{r^{2}}\right)}.$ (51)
@@
$L(P)=\frac{9k^{2}P\left(2\ 2^{3/4}m\sqrt[4]{-P}-\sqrt{q_{e}}\sqrt{4-\frac{2\sqrt{2}k^{2}\sqrt{-P}}{q_{e}}}\right)}{20q_{e}^{5/2}}+\frac{8\sqrt{2-\frac{\sqrt{2}k^{2}\sqrt{-P}}{q_{e}}}\left(k^{2}\sqrt{-P}+2\sqrt{2}q_{e}\right)}{15k^{2}q_{e}}.$ (52)
@@
Figure 3: Plot of the functions $F(r)$, $F(P)$, and $L(F)$ to the polymerized solution (45), with $q_{e}=k=0.5m$. The extreme points $P_{1}$ and $P_{2}$ inform us about the points where the electromagnetic Lagrangian changes its behavior.
@@ VI Energy Conditions
Subscript 1 refers to the case where the energy density is taken together with the radial pressure, while subscript 2 corresponds to the combination with the tangential pressure. Subscript 3, in turn, is used when only the energy density is considered, as in $WEC_{3}$, or when it is taken together with both pressures, as in $SEC_{3}$.
@@
The energy conditions for the Simpson-Visser case have already been analyzed for each field in the literature [Rodrigues:2023vtm]. Therefore, we now need to check for solutions inspired by LQG.
@@ VI.2 Polymerized solution
$DEC_{2}[\phi]=2WEC_{3}[\phi]\Longrightarrow\frac{28k^{2}m}{5r^{5}}-\frac{2\sqrt{1-\frac{k^{2}}{r^{2}}}\left(21k^{4}+8k^{2}r^{2}+16r^{4}\right)}{15k^{2}r^{4}}\geq 0,$ (75)
@@ VII Conclusion
The first model we studied was the Simpson-Visser spacetime. This model was chosen because it has been extensively studied in the literature and served as a consistency test for our method. The results obtained, both for magnetic and electric charges, differ from those presented in [Rodrigues:2023vtm, Alencar:2024yvh]. However, this difference arises because of the chosen coordinate system, so despite being different, our results are consistent with those found in the literature.
@@
Once the fields have been separated, we analyzed the energy conditions for each field individually. Since the Simpson-Visser case has been extensively studied in the literature, we did not address it here. For the second model, we found that all the energy conditions related to the electromagnetic field are always satisfied, whereas the scalar field can violate all the energy conditions, both inside and outside the event horizon, which is consistent with a phantom scalar field. For the third model, some of the analytical expressions are not very clear. However, we observe that, for points far from the event horizon, the inequality related to $NEC[E]$ is not satisfied, ensuring that, at least in some regions, the NEC is violated. Moreover, for the scalar field, it becomes more evident that the NEC is violated in all regions of spacetime. In summary, NEC violations are consistent with the behavior expected for BB solutions, such as those discussed here [Zaslavskii:2010qz].
The stability properties of solutions are known to depend critically on their source characteristics [Gonzalez:2008wd, Bronnikov:2013coa, Bronnikov:2021xao]. Consequently, classifying the source types capable of producing regular solutions is essential, which is the principal aim of this preliminary investigation. A natural continuation of our work is to assess viability through dynamical and observational probes, such as quasinormal modes, potential echoes, and long-lived modes in ultracompact regimes, in order to compare the stability of regular BHs and BB families sourced by different matter sectors. In particular, we will confront the sources obtained here with anisotropic fluids [Duran-Cabaces:2025sly, Franzin:2023slm].
It is well known that photons do not follow null geodesics in NED. As a consequence, photons may propagate with speeds larger or smaller than the vacuum speed of light. Causality in NED has been extensively investigated in the context of standard regular BHs and is also linked to the energy conditions [dePaula:2024yzy, Tomizawa:2023vir]. In future work, we intend to extend this causality analysis, and its connection to the energy conditions, to BB spacetimes supported by NED, as discussed here. We will also compare the resulting phenomenology in the effective and background metrics, focusing on light rings, shadows, and lensing [dePaula:2023ozi]. Taken together with shadow and lensing constraints, these tests will help delineate which sources support solutions that are regular, causal, and observationally consistent.
It is also possible to consider these spacetimes within alternative theories of gravity, such as $f(R)$ gravity or other modified gravity models, so that the matter sources are modified. This can mitigate violations of the energy conditions and reduce the need for phantom fields [GRAVMOD4, GRAVMOD3].